OpenStudy (anonymous):

Factor the GCF first and then factor the resulting binomial: 2x3 - 98x 2x(x + 49)(x - 49) 2x(x + 7)(x - 7) 2x(x + 14)(x - 14) prime

OpenStudy (anonymous):

GCF = greatest common factor what is the greatest common factor of 2x^3 and -98x?

OpenStudy (anonymous):

me thinks 2x is.... 2x(x^2 - 49) now factor x^2 - 49 x^2 - 49 = (x+7)(x-7) so we end up with 2x(x+7)(x-7)

OpenStudy (anonymous):

how did you get 49?

OpenStudy (anonymous):

when i factored out 2x, 98/2 = 49

OpenStudy (anonymous):

o okay thank you!

OpenStudy (anonymous):

you're welcome!